I need help on.....
on one of my homework problems i don't understand how to write a equation in slope-intercept form.... like for example one of my questions says
it shows a chart
Contains- Is perpendicular to.....
(3,-5) 5x+2y=10
i'm not asking for the answer, im just wondering what the steps are to solve the problem.
please e-mail me back soon
thanks

Answer by Earlsdon(6098)

Here are the steps for solving this problem:
1) The slope-intercept form of a linear equation is: y = mx + b where m is the slope of the line and b is the y-intercept (where the line crosses the y-axis) of the line. Get your equation into the slope-intercept form.
(y = -(5/2)x + 5)
2) Find the slope of your new equation. Hint: If your new line is perpendicular to another line, then the slope of your new line is the negative reciprocal of the slope of the given line. (The negative reciprocal of -(5/2) is (2/5))
3) Substitute into the slope-intercept form (y = mx + b) the value you find for your slope (m) and the x and y-coordinates of the given point (3, -5) and then solve this for the value of b, the y-intercept.
4) Write the final equation in the form y = mx + b but with your value of m (found in step 2)and your value of b (found in step 3).